# AnomalyDINO: Boosting Patch-based Few-shot Anomaly Detection with DINOv2

### v1

@@ 4.2 Ablation Study
We conduct additional experiments to assess the effect of specific design choices in our pipeline. The full results are provided in Appendix C. Here, we briefly describe the main insights.
@@
Comparing the two different preprocessing settings, ‘agnostic’ (our default) and ‘informed’ we observe that the agnostic preprocessing slightly outperforms the informed counterpart, see Figures 8 and 9. Depending on the product category suitable preprocessing steps lead to substantial improvements. In some cases augmenting the reference sample with rotations also negatively impacts the detection performance. For the full discussion, see Section C.1.
Inference time
Data augmentation increases the computational complexity of the proposed pipeline, as well as the memory bank size, and thus, the runtime of our nearest neighborhood approach. [^5] In practice, our implementation leverages GPU-accelerated neighborhood search (Johnson et al., 2019), leading to only slightly increased inference time. With increasing size of $\mathcal{M}$ reduction techniques such as coreset subsampling (Roth et al., 2022) are advisable. The effects of masking and augmentations on the inference time, presented in Figure 11, are not negligible but can be reduced by approximately 30% in the ‘informed’ scenario (on MVTec-AD).
Aggregation statistic
@@
Architecture size
In addition, we evaluated our pipeline with DINOv2 at different scales (ViT-S, ViT-B, ViT-L). The full results are given in Section C.3, Figure 12. We observe no considerable differences, in particular, also larger backbones give state-of-the-art results on MVTec-AD (all $k$) and VisA ($k=16$). Interestingly, larger architectures do not necessarily translate to higher performance. As the dimensionality of the feature space $\mathcal{F}$ increases with larger architectures, the distances between nominal patches also increases. The smallest model demonstrates the best performance on MVTec-AD, which primarily consists of single objects and less complex scenes. In contrast, the larger architecture size L performs better on VisA, which often involves multiple and more complex objects. This suggests the importance of achieving the ‘just right’ level of abstraction for the specific task, particularly in the few-shot regime.
@@ 5 Conclusion
This paper proposes a vision-only approach for one- and few-shot anomaly detection. Our method is based on similarities between patch representations extracted by DINOv2 (Oquab et al., 2024). We carefully design means to populate the nominal memory bank with diverse and relevant features utilizing zero-shot segmentation and simple data augmentation. Industrial settings require fast deployment, easy debugging and error correction, and rapid adaptation for covariance shifts in the normal data distribution. Our pipeline caters to these requirements through its simplicity and computational efficiency. The proposed method, AnomalyDINO, achieves state-of-the-art results on MVTec-AD and rivals all competing methods on VisA while outperforming all other training-free approaches. Its simplicity and strong performance render AnomalyDINO an excellent candidate for practitioners for industrial anomaly detection and an effective baseline for assessing few-shot and even full-shot anomaly detection in ongoing research.
@@
The specific pipeline exemplified in the paper focuses on simplicity and high throughput. However, the individual parts of our pipeline can easily be exchanged for more sophisticated alternatives. For example, our simple masking approach could be replaced by more specialized and adaptive masking techniques (which may also be relevant to other methods building on features extracted by DINOv2), and the simple upsampling and smoothing approach could be substituted by more sophisticated methods (e.g., Fu et al., 2024) to obtain the anomaly map from patch features. It would be interesting to see if that leads to further improvements in anomaly detection and localization, thereby further reducing the gap between few- and full-shot anomaly detectors. We also plan to improve the batched zero-shot performance of AnomalyDINO.
@@
Advancing few-shot visual anomaly detection methods can offer various benefits by enhancing manufacturing quality control through the rapid identification of defects with minimal nominal examples, which improves efficiency, reduces waste, and improves overall safety in the product lifecycle. Similar positive benefits can be expected outside the industrial domain, e.g., for healthcare diagnostics or environmental monitoring. It is, however, essential to recognize the shortcomings of automated anomaly detection systems. We believe that simpler methods can be adapted more quickly, monitored more effectively, and are therefore more reliable. In this context, awareness of the risks of overreliance must be heightened (see Appendix B for identified failure cases for the proposed method). In addition, strong visual anomaly detectors could also lead to potentially malicious or unintended uses. To address these concerns, including potential privacy infringements and possible socioeconomic impacts of automation, strategies such as establishing robust data governance, and implementing strict privacy protections are essential. Additionally, investing in workforce development can help manage the socioeconomic effects of automation and leverage the full potential of strong visual anomaly detection.
@@ Appendix A Detailed experimental results
In this section, we present the results of the default setting (AnomalyDINO-S and ‘agnostic’ preprocessing) in detail. The full results per category for MVTec-AD and VisA are given in Tables 5 and 6. Additional visualizations of anomaly maps are presented in Figures 3 and 4 for MVTec-AD and VisA, respectively.
@@
Figure 3: Examples – MVTec-AD. Depicted are, from left to right, a test sample per category (query), the ground truth anomaly annotation (GT), and the predicted anomaly map from AnomalyDINO in the 1- and 8-shot settings. The color coding is normalized by the maximum score on the ‘good’ test samples (per category).
Table 5: Detailed results on MVTec-AD. Reported are results on anomaly detection (image-level AUROC) and segmentation (PRO) of AnomalyDINO with default preprocessing (mean and standard deviation over three independent runs for each metric, all results in %).
@@
Bottle 99.5$\pm$0.4 96.1$\pm$0.3 99.7$\pm$0.1 96.4$\pm$0.1 99.9$\pm$0.1 96.6$\pm$0.2 100.0$\pm$0.0 96.6$\pm$0.1 100.0$\pm$0.0 96.5$\pm$0.2
Cable 91.6$\pm$1.2 88.2$\pm$0.7 92.1$\pm$0.7 88.5$\pm$0.7 93.8$\pm$1.4 89.5$\pm$0.5 94.7$\pm$0.3 89.6$\pm$0.3 94.4$\pm$1.1 89.5$\pm$0.7
Capsule 89.8$\pm$5.9 96.4$\pm$0.4 88.9$\pm$7.1 96.7$\pm$0.5 95.4$\pm$0.9 97.3$\pm$0.1 94.7$\pm$1.3 97.4$\pm$0.1 95.5$\pm$0.3 97.5$\pm$0.1
Carpet 100.0$\pm$0.0 97.9$\pm$0.1 100.0$\pm$0.0 97.9$\pm$0.0 100.0$\pm$0.0 97.7$\pm$0.1 100.0$\pm$0.0 97.7$\pm$0.1 100.0$\pm$0.0 97.6$\pm$0.2
Grid 99.2$\pm$0.1 96.1$\pm$0.0 99.2$\pm$0.1 96.3$\pm$0.1 99.2$\pm$0.1 96.2$\pm$0.2 99.4$\pm$0.2 96.2$\pm$0.1 99.8$\pm$0.3 95.9$\pm$0.3
Hazelnut 98.1$\pm$1.1 97.5$\pm$0.4 99.7$\pm$0.4 97.7$\pm$0.4 99.8$\pm$0.1 97.8$\pm$0.1 100.0$\pm$0.0 97.8$\pm$0.2 100.0$\pm$0.0 97.7$\pm$0.1
Leather 100.0$\pm$0.0 97.9$\pm$0.2 100.0$\pm$0.0 97.5$\pm$0.3 100.0$\pm$0.0 97.1$\pm$0.3 100.0$\pm$0.0 97.0$\pm$0.1 100.0$\pm$0.0 96.7$\pm$0.1
Metal nut 99.9$\pm$0.2 93.8$\pm$0.1 100.0$\pm$0.1 94.4$\pm$0.3 100.0$\pm$0.0 95.1$\pm$0.1 100.0$\pm$0.0 95.2$\pm$0.3 100.0$\pm$0.0 95.5$\pm$0.1
Pill 93.1$\pm$1.5 96.9$\pm$0.2 94.9$\pm$2.2 97.1$\pm$0.2 96.3$\pm$0.1 97.3$\pm$0.0 97.1$\pm$0.4 97.4$\pm$0.1 97.6$\pm$0.3 97.5$\pm$0.1
Screw 90.1$\pm$1.0 92.5$\pm$0.2 92.3$\pm$0.2 93.5$\pm$0.0 91.3$\pm$2.4 93.6$\pm$0.7 92.5$\pm$0.8 94.6$\pm$0.4 93.5$\pm$1.4 95.3$\pm$0.2
Tile 100.0$\pm$0.0 88.6$\pm$0.2 100.0$\pm$0.0 88.2$\pm$0.2 100.0$\pm$0.0 87.9$\pm$0.2 100.0$\pm$0.0 87.5$\pm$0.2 100.0$\pm$0.0 86.9$\pm$0.6
Toothbrush 99.4$\pm$0.5 94.5$\pm$0.1 99.9$\pm$0.1 94.9$\pm$0.3 99.9$\pm$0.1 95.0$\pm$0.3 99.9$\pm$0.1 95.5$\pm$0.5 99.6$\pm$0.3 95.9$\pm$0.6
Transistor 90.9$\pm$2.1 64.4$\pm$0.5 89.2$\pm$5.9 64.8$\pm$3.1 91.9$\pm$3.5 66.5$\pm$1.4 96.4$\pm$1.8 70.3$\pm$1.6 98.3$\pm$0.5 72.7$\pm$1.0
Wood 98.8$\pm$0.4 93.9$\pm$0.3 98.7$\pm$0.6 93.8$\pm$0.2 98.6$\pm$0.6 93.8$\pm$0.1 98.6$\pm$0.2 93.8$\pm$0.2 98.2$\pm$0.4 93.5$\pm$0.5
Zipper 98.9$\pm$0.1 83.2$\pm$0.4 99.0$\pm$0.1 84.4$\pm$0.7 99.3$\pm$0.2 85.6$\pm$0.2 99.2$\pm$0.2 85.5$\pm$0.6 99.3$\pm$0.3 85.9$\pm$0.7
Mean 96.6$\pm$0.5 91.9$\pm$0.1 96.9$\pm$0.8 92.1$\pm$0.2 97.7$\pm$0.1 92.5$\pm$0.1 98.2$\pm$0.1 92.8$\pm$0.0 98.4$\pm$0.1 93.0$\pm$0.1
query GT 1-shot 8-shot query GT 1-shot 8-shot
Figure 4: Examples – VisA. Depicted are, from left to right, a test sample per category (query), the ground truth anomaly annotation (GT), and the predicted anomaly map from AnomalyDINO in the 1- and 8-shot settings. The color bar is normalized by the maximum score on the ‘good’ test samples (per category). For the category ‘Capsules’ (left column, second from top) we changed the color map for better visibility. Best viewed at a higher zoom level as some anomalies are quite small.
Table 6: Detailed results on VisA. Reported are results on anomaly detection (image-level AUROC) and segmentation (PRO) of AnomalyDINO with default preprocessing (mean and standard deviation over three independent runs for each metric, all results in %).
@@
Candle 83.8$\pm$0.7 77.6$\pm$0.6 86.2$\pm$3.8 78.4$\pm$0.1 88.0$\pm$3.0 79.3$\pm$0.3 89.9$\pm$1.0 79.7$\pm$0.2 92.0$\pm$0.8 80.7$\pm$0.3
Capsules 98.4$\pm$0.3 80.4$\pm$0.5 99.1$\pm$0.2 81.0$\pm$0.4 99.1$\pm$0.2 81.5$\pm$0.2 99.1$\pm$0.5 81.9$\pm$0.1 98.8$\pm$0.4 82.4$\pm$0.2
Cashew 85.3$\pm$5.6 95.0$\pm$1.5 89.3$\pm$5.6 95.7$\pm$1.1 93.8$\pm$0.5 96.2$\pm$0.8 95.1$\pm$0.7 95.9$\pm$0.3 95.5$\pm$0.1 96.2$\pm$0.3
Chewinggum 98.2$\pm$0.3 90.6$\pm$0.8 98.7$\pm$0.3 91.8$\pm$0.8 98.8$\pm$0.2 92.2$\pm$0.5 98.9$\pm$0.1 92.5$\pm$0.3 99.0$\pm$0.0 92.6$\pm$0.0
Fryum 95.1$\pm$0.3 92.6$\pm$0.3 96.6$\pm$0.5 93.2$\pm$0.1 97.2$\pm$0.5 93.6$\pm$0.1 97.9$\pm$0.6 93.8$\pm$0.1 97.8$\pm$0.6 94.0$\pm$0.1
Macaroni1 83.4$\pm$3.0 86.7$\pm$1.0 83.3$\pm$1.7 87.6$\pm$0.2 85.8$\pm$0.9 88.7$\pm$0.2 86.9$\pm$0.8 89.6$\pm$0.2 88.2$\pm$0.7 90.0$\pm$0.2
Macaroni2 61.4$\pm$3.8 66.3$\pm$0.7 67.1$\pm$5.3 67.5$\pm$0.3 70.0$\pm$1.8 68.5$\pm$0.4 74.0$\pm$1.2 70.3$\pm$0.6 76.5$\pm$1.5 71.9$\pm$0.5
PCB1 89.0$\pm$1.2 90.6$\pm$0.3 88.9$\pm$1.5 90.6$\pm$0.6 92.3$\pm$2.3 91.2$\pm$0.4 92.9$\pm$0.6 91.8$\pm$0.2 95.2$\pm$0.8 92.5$\pm$0.3
PCB2 87.8$\pm$2.0 86.5$\pm$1.1 89.4$\pm$2.5 87.5$\pm$0.2 91.8$\pm$0.7 88.4$\pm$0.1 92.8$\pm$0.6 89.1$\pm$0.2 93.0$\pm$0.4 89.5$\pm$0.3
PCB3 84.1$\pm$2.4 85.6$\pm$0.8 88.5$\pm$1.3 87.1$\pm$0.6 91.8$\pm$0.3 87.8$\pm$0.4 93.8$\pm$0.7 89.4$\pm$0.6 94.5$\pm$0.2 90.0$\pm$0.2
PCB4 76.7$\pm$11.5 79.8$\pm$5.0 86.5$\pm$7.2 83.3$\pm$3.7 94.9$\pm$3.0 86.1$\pm$0.8 97.3$\pm$0.6 88.8$\pm$1.1 99.0$\pm$0.3 90.6$\pm$0.9
Pipe fryum 91.8$\pm$2.6 97.4$\pm$0.3 93.5$\pm$3.0 97.5$\pm$0.3 94.8$\pm$2.7 97.2$\pm$0.1 96.5$\pm$1.6 97.3$\pm$0.1 98.0$\pm$0.3 97.3$\pm$0.1
Mean 86.2$\pm$1.4 85.8$\pm$0.6 88.9$\pm$1.9 86.8$\pm$0.4 91.5$\pm$0.7 87.6$\pm$0.2 92.9$\pm$0.1 88.3$\pm$0.2 94.0$\pm$0.0 89.0$\pm$0.2
@@
In this context, we observe a peculiarity of the few-shot regime (which does not occur for sufficiently populated $\mathcal{M}$). The results crucially depend on the chosen reference image(s). The high variances, predominantly in the 1- and 2-shot setting, for category ‘Capsule’ for MVTec-AD, or ‘Cashew’ or ‘PCB4’ for VisA demonstrate this. We discuss the potential failure cases of choosing a sub-optimal reference sample in the following section (Section B.2).

### v2

@@ 4.2 Ablation Study
We conduct additional experiments to assess the effect of specific design choices in our pipeline. The full ablation study is provided in Appendix C, here we briefly summarize the main insights. Figure 3 supports the ablation study by highlighting the two key aspects: performance and runtime.
Figure 3: Detection AUROC vs. inference time per sample on MVTec-AD in the 1-shot setting. The input resolution is given in parentheses after the method name. All runtimes are measured on a single NVIDIA A40 if not stated otherwise. (Note that for ADP and WinCLIP+ no official code is available.)
Inference time
Comparing the inference times of AnomalyDINO with the other approaches in the 1-shot setting (see Figure 3), we observe that AnomalyDINO achieves significantly faster inference times than SOTA few-shot competitors (note the logarithmic scale). For example, AnomalyDINO-S takes roughly 60ms to process an image at a resolution of 448. The only approaches with lower inference time are PatchCore and AnomalyDINO with ViT-backbones trained on ImageNet, which however sacrifice performance. Notably, while augmentations increase the memory bank, and thus, the potential runtime of the nearest neighborhood search, we find that the effect is negligible in practice. [^5] Our implementation leverages GPU-accelerated neighborhood search [19], leading to only slightly increased inference time. With increasing size of $\mathcal{M}$ reduction techniques such as coreset subsampling [33] are advisable. A more detailed runtime analysis is included in Appendix C, Table 9.
@@
To study the impact of preprocessing, we compare the resulting detection AUROCs of AnomalyDINO with and without preprocessing in Figure 3. Across four different configurations of AnomalyDINO, we observe that by incorporating the proposed preprocessing, the resulting AUROCs increase by approximately $2\%$ without significantly affecting the inference time. The effect of proper preprocessing seems to increase with higher resolutions. Additionally, we compare the two different preprocessing settings, ‘agnostic’ (our default) and ‘informed’, and find that the agnostic preprocessing slightly outperforms the informed counterpart, see Figures 9 and 10. Depending on the product category, suitable preprocessing steps lead to substantial improvements. And while in principle augmenting the reference sample with rotations may negatively impact the detection performance, we observe this in only very few examples. For the full discussion, see Section C.1.
Aggregation choice
@@
Architecture size/choice
We also evaluated the proposed framework using Vision Transformers pre-trained on ImageNet (see Figure 3). While these backbones slightly outperform PatchCore, they give weaker features compared to DINOv2 and are incompatible with the proposed masking procedure (zero-shot masking does not function effectively). This underlines that DINOv2 is indeed excellently suited for visual AD. In addition, ViT-based architectures like DINOv2 easily allow handling images with varying resolutions. In this context, we find that operating at a resolution of 448 gives the best trade-off between performance and inference time, but even higher detection AUROCs are possible at higher resolutions.
In addition, we evaluated our pipeline with DINOv2 at different scales (ViT-S, ViT-B, ViT-L). The full results are given in Section C.3, Figure 14. We observe no considerable differences, in particular, also larger backbones give state-of-the-art results on MVTec-AD (all $k$) and VisA ($k\geq 8$). Interestingly, larger architectures do not necessarily translate to higher performance. The smallest model demonstrates the best performance on MVTec-AD, which primarily consists of single objects and less complex scenes. In contrast, the larger architecture sizes perform better on VisA, which often involves multiple and more complex objects. This suggests the importance of achieving the ‘just right’ level of abstraction for the specific task, particularly in the few-shot regime.
@@ 5 Conclusion
This paper proposes a vision-only approach for one- and few-shot anomaly detection. Our method is based on similarities between patch representations extracted by DINOv2 [29]. We carefully design means to populate the nominal memory bank with diverse and relevant features while minimizing false positives utilizing zero-shot segmentation and simple data augmentation. Industrial settings require fast deployment, easy debugging and error correction, and rapid adaptation for covariance shifts in the normal data distribution. Our pipeline caters to these requirements through its simplicity and computational efficiency. The proposed method, AnomalyDINO, achieves state-of-the-art results on MVTec-AD and rivals all competing methods on VisA while outperforming all other training-free approaches. Thus, our approach achieves the best of both worlds: improved performance and reduced inference time, especially compared to more complex vision-language methods. Its simplicity and strong performance render AnomalyDINO an excellent candidate for practitioners for industrial anomaly detection and an effective baseline for assessing few-shot and even full-shot anomaly detection in ongoing research.
@@
The specific pipeline exemplified in the paper focuses on simplicity and high throughput. However, the individual parts of our pipeline can easily be exchanged for more sophisticated alternatives. For example, our simple masking approach could be replaced by more specialized and adaptive masking techniques (which may also be relevant to other methods building on features extracted by DINOv2), and the simple upsampling and smoothing approach could be substituted by more sophisticated methods like [14] to obtain the anomaly map from patch features. It would be interesting to see if that leads to further improvements in anomaly detection and localization, thereby further reducing the gap between few- and full-shot anomaly detectors. We also plan to improve the batched zero-shot performance of AnomalyDINO.
@@
Advancing few-shot visual anomaly detection methods can offer various benefits by enhancing manufacturing quality control through the rapid identification of defects with minimal nominal examples, which improves efficiency, reduces waste, and improves overall safety in the product lifecycle. Similar positive benefits can be expected outside the industrial domain, e.g., for healthcare diagnostics or environmental monitoring. It is, however, essential to recognize the shortcomings of automated anomaly detection systems. We believe that simpler methods can be adapted more quickly, monitored more effectively, and are therefore more reliable. In this context, awareness of the risks of overreliance must be heightened (see Appendix B for identified failure cases of the proposed method). In addition, strong visual anomaly detectors could also lead to potentially malicious or unintended uses. To address these concerns, including potential privacy infringements and possible socioeconomic impacts of automation, strategies such as establishing robust data governance, and implementing strict privacy protections are essential. Additionally, investing in workforce development can help manage the socioeconomic effects of automation and leverage the full potential of strong visual anomaly detection.
@@ Appendix A Detailed experimental results
In this section, we present additional visualization and results for the proposed method.
@@
Figure 4: Examples – MVTec-AD. Depicted are, from left to right, a test sample per category (query), the ground truth anomaly annotation (GT), and the predicted anomaly map from AnomalyDINO-S (448) in the 1- and 8-shot settings. The color coding is normalized by the max. score over ‘good’ test samples.
Further anomaly maps, predicted by AnomalyDINO, are presented in Figures 4 and 5. The full results per category for MVTec-AD and VisA are given in Tables 5 and 6, respectively.
Table 5: Detailed results on MVTec-AD. Reported are results on anomaly detection (image-level AUROC) and segmentation (PRO) of AnomalyDINO-S (672) with default preprocessing (mean and standard deviation over three independent runs, all results in %).
@@
Bottle 99.7$\pm$0.2 95.9$\pm$0.3 99.9$\pm$0.1 96.3$\pm$0.0 99.9$\pm$0.1 96.7$\pm$0.1 100.0$\pm$0.0 96.5$\pm$0.4 99.9$\pm$0.1 96.4$\pm$0.4
Cable 92.7$\pm$0.8 89.4$\pm$0.4 92.4$\pm$1.1 89.5$\pm$0.4 93.8$\pm$0.9 90.4$\pm$0.3 95.2$\pm$0.3 90.5$\pm$0.2 95.1$\pm$0.6 90.5$\pm$0.6
Capsule 90.2$\pm$5.5 97.1$\pm$0.3 89.2$\pm$7.9 97.3$\pm$0.6 95.8$\pm$0.5 97.9$\pm$0.1 95.6$\pm$0.5 97.9$\pm$0.1 95.5$\pm$0.6 98.1$\pm$0.1
Carpet 100.0$\pm$0.0 97.8$\pm$0.0 100.0$\pm$0.0 97.9$\pm$0.0 100.0$\pm$0.0 97.8$\pm$0.0 100.0$\pm$0.0 97.8$\pm$0.0 100.0$\pm$0.0 97.8$\pm$0.0
Grid 99.1$\pm$0.2 97.2$\pm$0.1 99.2$\pm$0.4 97.2$\pm$0.1 99.5$\pm$0.3 97.2$\pm$0.1 99.5$\pm$0.1 97.2$\pm$0.0 99.7$\pm$0.3 97.0$\pm$0.2
Hazelnut 97.5$\pm$2.6 97.4$\pm$0.4 99.6$\pm$0.5 98.0$\pm$0.3 99.8$\pm$0.1 98.0$\pm$0.1 100.0$\pm$0.0 98.1$\pm$0.1 100.0$\pm$0.0 98.1$\pm$0.1
Leather 100.0$\pm$0.0 97.9$\pm$0.1 100.0$\pm$0.0 97.8$\pm$0.0 100.0$\pm$0.0 97.6$\pm$0.1 100.0$\pm$0.0 97.6$\pm$0.1 100.0$\pm$0.0 97.3$\pm$0.2
Metal nut 99.9$\pm$0.1 94.2$\pm$0.0 100.0$\pm$0.0 94.6$\pm$0.2 100.0$\pm$0.0 95.3$\pm$0.1 100.0$\pm$0.0 95.4$\pm$0.3 100.0$\pm$0.0 95.7$\pm$0.1
Pill 93.7$\pm$0.9 97.3$\pm$0.1 95.4$\pm$0.7 97.5$\pm$0.1 96.0$\pm$0.2 97.6$\pm$0.1 97.2$\pm$0.2 97.7$\pm$0.1 97.9$\pm$0.1 97.8$\pm$0.1
Screw 93.2$\pm$0.3 93.4$\pm$0.4 93.5$\pm$0.8 94.3$\pm$0.4 92.7$\pm$2.3 94.5$\pm$0.9 93.5$\pm$1.1 95.2$\pm$0.5 94.7$\pm$0.9 95.9$\pm$0.3
Tile 100.0$\pm$0.0 88.0$\pm$0.2 100.0$\pm$0.0 87.6$\pm$0.4 100.0$\pm$0.0 87.1$\pm$0.4 100.0$\pm$0.0 86.7$\pm$0.2 100.0$\pm$0.0 86.0$\pm$0.5
Toothbrush 97.4$\pm$0.5 94.0$\pm$0.8 98.1$\pm$1.0 94.7$\pm$0.3 97.5$\pm$0.6 94.7$\pm$0.3 97.7$\pm$1.6 95.1$\pm$0.8 98.1$\pm$1.8 95.8$\pm$1.0
Transistor 90.9$\pm$1.2 67.3$\pm$2.1 89.4$\pm$4.6 68.4$\pm$3.1 93.2$\pm$2.2 70.6$\pm$1.4 96.2$\pm$1.3 75.3$\pm$1.9 97.6$\pm$0.3 78.2$\pm$1.4
Wood 98.0$\pm$0.2 94.7$\pm$0.1 98.0$\pm$0.1 94.6$\pm$0.0 97.9$\pm$0.2 94.6$\pm$0.1 98.3$\pm$0.4 94.4$\pm$0.3 98.3$\pm$0.6 94.2$\pm$0.4
Zipper 97.4$\pm$0.9 89.2$\pm$1.2 98.9$\pm$0.4 90.2$\pm$0.4 99.0$\pm$0.4 91.2$\pm$0.3 99.6$\pm$0.1 91.1$\pm$0.4 99.6$\pm$0.3 91.7$\pm$0.5
Mean 96.6$\pm$0.4 92.7$\pm$0.1 96.9$\pm$0.7 93.1$\pm$0.2 97.7$\pm$0.2 93.4$\pm$0.1 98.2$\pm$0.2 93.8$\pm$0.1 98.4$\pm$0.1 94.0$\pm$0.1
Table 6: Detailed results on VisA. Reported are results on anomaly detection (image-level AUROC) and segmentation (PRO) of AnomalyDINO-S (672) with default preprocessing (mean and standard deviation over three independent runs, all results in %).
@@
Candle 87.9$\pm$0.3 96.8$\pm$0.4 89.4$\pm$3.0 97.0$\pm$0.2 91.3$\pm$2.9 97.2$\pm$0.1 93.5$\pm$1.2 97.3$\pm$0.2 94.5$\pm$0.5 97.6$\pm$0.2
Capsules 98.4$\pm$0.5 95.1$\pm$0.7 98.9$\pm$0.1 95.5$\pm$0.2 99.2$\pm$0.1 96.3$\pm$0.4 99.2$\pm$0.1 96.7$\pm$0.2 99.2$\pm$0.2 97.2$\pm$0.3
Cashew 86.1$\pm$3.6 96.1$\pm$0.9 89.4$\pm$3.8 96.7$\pm$0.7 94.5$\pm$0.7 97.4$\pm$0.5 95.3$\pm$0.6 97.3$\pm$0.2 96.0$\pm$0.3 97.3$\pm$0.1
Chewinggum 98.0$\pm$0.4 92.0$\pm$1.0 98.6$\pm$0.4 92.9$\pm$0.3 98.8$\pm$0.2 93.0$\pm$0.1 98.8$\pm$0.2 93.1$\pm$0.3 98.8$\pm$0.2 93.0$\pm$0.3
Fryum 94.8$\pm$0.5 93.2$\pm$0.2 96.5$\pm$0.2 93.9$\pm$0.3 97.0$\pm$0.1 94.5$\pm$0.4 97.6$\pm$0.4 94.9$\pm$0.3 97.9$\pm$0.2 95.1$\pm$0.0
Macaroni1 87.5$\pm$1.1 97.5$\pm$0.3 87.5$\pm$0.9 97.9$\pm$0.3 89.5$\pm$1.4 98.3$\pm$0.2 90.1$\pm$1.7 98.6$\pm$0.2 90.4$\pm$1.1 98.7$\pm$0.1
Macaroni2 62.2$\pm$4.3 92.0$\pm$0.7 66.9$\pm$1.9 93.0$\pm$0.4 70.0$\pm$1.7 93.9$\pm$0.8 74.9$\pm$0.4 95.0$\pm$0.6 77.6$\pm$0.8 95.7$\pm$0.3
PCB1 91.5$\pm$2.0 92.6$\pm$0.2 91.2$\pm$2.7 92.5$\pm$0.5 94.0$\pm$2.1 93.3$\pm$0.5 95.5$\pm$0.5 93.9$\pm$0.2 96.8$\pm$0.7 94.2$\pm$0.2
PCB2 84.8$\pm$1.2 89.9$\pm$0.2 88.1$\pm$2.5 90.7$\pm$0.3 91.1$\pm$1.7 91.4$\pm$0.2 92.6$\pm$0.3 92.0$\pm$0.1 93.2$\pm$0.1 92.5$\pm$0.2
PCB3 84.9$\pm$3.3 88.5$\pm$1.3 89.4$\pm$3.8 90.8$\pm$0.5 94.3$\pm$0.4 91.7$\pm$0.4 95.6$\pm$0.2 93.1$\pm$0.3 96.5$\pm$0.3 93.9$\pm$0.3
PCB4 79.9$\pm$13.7 78.5$\pm$6.8 87.4$\pm$11.3 82.0$\pm$6.1 96.2$\pm$2.6 84.1$\pm$1.5 98.0$\pm$0.3 87.9$\pm$2.3 99.0$\pm$0.4 90.4$\pm$1.8
Pipe fryum 92.7$\pm$2.7 98.0$\pm$0.0 93.3$\pm$1.2 97.9$\pm$0.2 94.6$\pm$1.9 97.8$\pm$0.1 95.0$\pm$1.6 97.6$\pm$0.1 97.2$\pm$0.9 97.7$\pm$0.1
Mean 87.4$\pm$1.2 92.5$\pm$0.5 89.7$\pm$1.3 93.4$\pm$0.6 92.6$\pm$0.9 94.1$\pm$0.1 93.8$\pm$0.3 94.8$\pm$0.2 94.8$\pm$0.2 95.3$\pm$0.2
query GT 1-shot 8-shot query GT 1-shot 8-shot
Figure 5: Examples – VisA (1/2). Depicted are, from left to right, a test sample per category (query), the ground truth anomaly annotation (GT), and the predicted anomaly map from AnomalyDINO-S (448) in the 1- and 8-shot settings. The color bar is normalized by the maximum score on the ‘good’ test samples (per category). For the category ‘Capsules’ (left column, second from top) we changed the color map for better visibility. Best viewed at a higher zoom level as some anomalies are quite small.
@@
In this context, we observe a peculiarity of the few-shot regime (which does not occur for sufficiently populated $\mathcal{M}$). The results crucially depend on the chosen reference image(s).
Table 7: Full-shot results on MVTec-AD and VisA with AnomalyDINO-S in the default setting (no std reported as results are deterministic when all samples are considered).
Dataset Resolution Detection Segmentation
AUROC F1-max AP AUROC F1-max PRO
MVTec-AD 448 99.3 98.8 99.7 97.9 61.8 93.9
672 99.5 99.0 99.8 98.2 64.3 95.0
VisA 448 97.2 93.7 97.6 98.7 50.5 95.0
672 97.6 94.5 98.0 98.8 53.8 96.1
The high variances, predominantly in the 1- and 2-shot setting, for category ‘Capsule’ for MVTec-AD, or ‘Cashew’ or ‘PCB4’ for VisA demonstrate this. We discuss the potential failure cases of choosing a sub-optimal reference sample in the following section (Section B.2).
Finally, we also report the full-shot performance, see Table 7. As the diversity within $\mathcal{M}$ is sufficiently high given the large number of reference samples, we only apply masing here (no augmentations). The results demonstrate that the performance nicely improves for all considered metrics.